OpenStudy (anonymous):

Jerry invested $7000, part at 8% simple interest and the rest at 5% simple interest for a period of 1 year. If he received a total annual interest of $476 from both investments, how much did he invest at each rate? I am bad with word problems. can someone please help me and explain to how to do this equation. Thanks.

OpenStudy (anonymous):

let jerry invest x and 7000-x in 8% and 5% respectively thus the interest he recieved after 1 yr are 8x/100 and 5(7000-x)/100 as per the quest. we have 8x/100 + 5(7000-x)/100=476 or 8x+35000-5x=47600 or 3x =47600-35000 or x= 12600/3 or x= 4200 hence Jerry invested 4200 in 8% and (7000-4200)=2800 in 5% one

OpenStudy (anonymous):

where did the 100 come from. sorry am trying to understand where you got that from.